# Computation of $\left\langle\Phi^{2}\right\rangle$ and quantum fluxes at the polar interior of a spinning black hole

### v1

@@ I Introduction
In the case that the quantity of interest is the field square $\Phi^{2}$ or the stress-energy tensor $T_{\mu\nu}$, we respectively denote the unrenormalized bi-tensor by $\frac{1}{2}G_{\Psi}(x,x^{\prime})$ or $\left\langle T_{\mu\nu}(x,x^{\prime})\right\rangle^{\Psi}$, the counterterm by $\frac{1}{2}G^{\text{CT}}(x,x^{\prime})$ or $T_{\mu\nu}^{\text{CT}}(x,x^{\prime})$, and the point-splitting regularization procedure then amounts to $\left\langle\Phi^{2}(x)\right\rangle_{\text{ren}}^{\Psi}=\frac{1}{2}\lim_{x^{\prime}\to x}\left(G_{\Psi}(x,x^{\prime})-G^{\text{CT}}(x,x^{\prime})\right)$ or $\left\langle T_{\mu\nu}(x)\right\rangle_{\text{ren}}^{\Psi}=\lim_{x^{\prime}\to x}\left(\left\langle T_{\mu\nu}(x,x^{\prime})\right\rangle^{\Psi}-T_{\mu\nu}^{\text{CT}}(x,x^{\prime})\right)$. In this paper, for the two-point function $G_{\Psi}(x,x^{\prime})$ we shall later make the choice of the anticommutator $G_{\Psi}^{(1)}(x,x^{\prime})\equiv\left\langle\left\{\Phi(x),\Phi(x^{\prime})\right\}\right\rangle_{\Psi}$, also called the Hadamard two-point function (HTPF), although other choices for $G_{\Psi}(x,x^{\prime})$ are also possible.
@@
The literature results mentioned so far were for spherically-symmetric BHs. In the astrophysically most important case of a stationary, rotating (Kerr) BH, because of the lack of spherical symmetry, $\theta$-splitting is not possible and, because of the lack of staticity, the Euclideanization technique is in principle not implementable either. Thus, progress in Kerr was for a long time made only by calculating differences of renormalized expectation values of quantities in two different states. This was some times combined with the expected behavior of the RSET for one of the states in some specific spacetime region (namely, at infinity or near a horizon) in order to apply the state subtraction method so as to gain knowledge about the behavior of the RSET for the other state in that region: see Refs. [41, 10, 43, 42] outside the EH and Ref. [2] on the CH (and [6] on the CH of Kerr-dS). An exception to that is the axis of symmetry of Kerr, where no rotation is ‘felt’, and that was used to directly obtain the RSET in a ‘formal’ Hartle-Hawking state on the pole of the EH in Refs. [55, 11]. An important technical breakthrough in renormalization was achieved using the $t$- and $\varphi$-splitting variants of PMR in [53], where the authors managed to calculate the RSET outside the EH of Kerr. This is so far the only time that the calculation of an RSET has been carried out outside a Kerr BH.
@@ II.4 Field quantization and Unruh state
In this subsection we sketch the quantization of the field and the definition of our quantum state of interest, namely the Unruh state. For details, we refer the reader to Secs. III and IV in Ref. [3]. The scalar field $\Phi$ is typically quantized by expanding it in terms of a choice of modes and then promoting the coefficients in the mode expansion to (creation and annihilation) operators. In the previous subsection we introduce the Eddington modes since they are convenient for practical calculations. However, the Unruh state is instead more naturally defined in terms of some other modes, the so-called Unruh modes, which consist of the union of the following two subsets of modes. The first subset, which merges with the Eddington $f^{\text{in}}_{\omega lm}$ in the BH exterior, is defined as having no upward excitations coming from $\{U\in\mathbb{R},V=0\}$ (i.e., from the union of $H_{L}$ and $H_{\text{past}}$, see Fig. 1) and having positive frequencies with respect to the Eddington coordinate $v$ along $\{U\to-\infty,V>0\}$ (i.e., along PNI, see Fig. 1); the second subset is defined by being positive frequency with respect to the Kruskal coordinate $U$ along $H_{L}\cup H_{\text{past}}$ and having no upward excitations from PNI. The Unruh state [82] is then defined as the quantum state which is annihilated by the annihilation operator coefficients when expanding the quantum field in Unruh modes. The Unruh state is the state of relevance for astrophysical BHs since it models a BH evaporating via the emission of quantum, thermal (Hawking) radiation.
@@ Appendix A Large-$l$ asymptotics
In this Appendix we briefly describe the large-$l$ analysis of the interior radial function $\psi_{\omega l}^{\text{int}}$ (see Eq. (30)) and the reflection coefficient $\rho^{\text{up}}_{\omega l}$ (see Eq. (27)) in a Kerr BH, focusing on the $m=0$ case, given that only $m=0$ modes are relevant at the pole. The analysis we describe yields the leading order expressions which are subsequently used in Sec. III.2.1 for the ID computation of $\left\langle\Phi^{2}\right\rangle$.
@@ B.1 Basic ID analysis
Thus, by taking minus the second derivative with respect to $\varepsilon$ of Eq. (109), we obtain the following distributional identity valid for $\sin\delta\neq\varepsilon$:  [^34] In this footnote we present an alternative derivation of Eq. (114). Strictly speaking, this integral is not well defined because the $\omega^{2}$ factor [or the $\omega$ factor in the analogous situation of the integral appearing in Eq. (126)] spoils convergence at large $\omega$. Nevertheless, the integral (like some other integrals involved in our $t$-splitting method) becomes well-defined with the procedure of “oscillation cancellation” (see Ref. [47]), and it is this “generalized integral” that we want to compute here. An alternative (and equivalent) formulation of this generalized integral is via multiplying the integrand by a regulating function $\exp(-c\,\omega)$ (with $c>0$) and then taking the limit $c\to 0$ after integration. This alternative definition, involving the $\exp(-c\,\omega)$ factor, is more commonly used in the mathematical physics literature; however, it is much harder to numerically implement it (compared to oscillation cancellation). Nevertheless, for the analytical derivation of this (generalized) integral – which is our concern here – this second method [i.e. using the regulating function $\exp(-c\,\omega)$] is very convenient. Thus, the relevant integrand now becomes $\omega^{2}\cos\left(\omega\varepsilon\right)J_{0}\left(\omega\sin\delta\right)\exp(-c\,\omega)$. Unfortunately, Mathematica is unable to directly perform this integral. However, when reduced to the evaluation of $\Re\lim_{c\to 0}\int_{0}^{\infty}\omega^{2}\exp\left(i\omega\varepsilon\right)J_{0}\left(\omega\sin\delta\right)\exp(-c\,\omega)$, Mathematica successfully yields the desired result given in Eq. (114).
@@
Finally, we deal with $P_{\text{div}}^{\left(2\right)}$. We denote:
$\Sigma_{2}\left(\delta,\omega\right)\equiv\sum_{l=0}^{\infty}\left(\lambda_{\omega l}\frac{S_{\omega l}\left(\delta\right)}{S_{\omega l}\left(0\right)}\right)$ (117)
@@ Appendix C Numerical methods: the computational side
This section deals with the numerical implementation of the $t$-splitting regularization procedure in practice — from the computation of the basic ingredients $\rho_{\omega l}^{\text{up}}$, $\psi_{\omega l}^{\text{int}}$ and $\psi_{\omega l,r}^{\text{int}}$ (as well as $S_{\omega l}(\theta=0)$) to the performance of the various steps (including various challenges that arise), leading to the renormalized quantities $\left\langle\Phi^{2}\right\rangle_{\text{ren}}^{U}$ and $\left\langle T_{yy}\right\rangle_{\text{ren}}^{U}$ in the cases under consideration (detailed in Sec. IV).
@@
Finally, the spheroidal eigenfunctions $S_{\omega l}(\theta=0)$ [defined via Eqs. (44), (20) and (19)] were computed using Mathematica’s built-in function SpheroidalPS. In particular, taking normalization into account, our $S_{\omega l}(\theta=0)$ corresponds to $\sqrt{(2l+1)/2}\;\text{SpheroidalPS}[l,0,i\,a\,\omega,1]$. The numerical evaluation of the spheroidal eigenfunctions was done to 300 significant digits.
@@ C.2 Numerical methods
First, in principle, the ID needs to be subtracted from the bare mode contribution, to be followed by a summation over $l$ (per $\omega$) to produce the basic integrand function [defined in Eq. (40)]. For $\Phi^{2}$ this ID subtraction stage (demonstrated in Fig. 3) merely includes the removal of the analytically-known $O\left(l^{0}\right)$ leading order $c_{0}$ [see Eq. (59)], leaving a remainder that decays like $1/l\left(l+1\right)$, which is to be summed over. As already mentioned above, we overcome this slow decay by fitting the sequence of partial sums (namely $\sum_{l^{\prime}=0}^{l}\left(E_{\omega l^{\prime}}-E_{\omega l^{\prime}}^{\text{div}}\right)$) with $\sim 100$ (or more) orders in $1/l$ (starting with order $l^{0}$). The fitted coefficient of this $l^{0}$ term then constitutes the desired $l$-sum. This allows the basic integrand $E^{\text{basic}}\left(\omega\right)$ to be extracted to a precision of typically $\gtrsim 50$ digits, depending on $r$ (the $l$-sum error is in fact evaluated by comparing fits with different numbers of orders in $1/l$).
@@
Hence, we now focus on the direct computation of $\left\langle T_{uu}\right\rangle_{\text{ren}}^{U}$. For $T_{uu}$, the large-$l$ divergent piece $T_{uu\left(\omega l\right)}^{\text{div}}$ has an analytically-known leading-order piece [$c_{2}l\left(l+1\right)$, with $c_{2}$ given in Eq. (74)] and an additional $O\left(l^{0}\right)$ piece [namely $c_{01}\left(\omega\right)$ defined in Eq. (76) ]. This quantity $c_{01}\left(\omega\right)$ is, in principle, not known analytically, but can be extracted numerically. Thus, the straightforward procedure would be to compute the $l$-sum in two stages: first, numerically extracting $c_{01}\left(\omega\right)$, and using it to obtain the regularized sequence $T_{uu\left(\omega l\right)}-T_{uu\left(\omega l\right)}^{\text{div}}$; and then summing this regularized sequence over $l$ to obtain $T_{uu}^{\text{basic}}\left(\omega\right)$. This procedure is demonstrated in Fig. 4. However, we find it simpler and more efficient (and significantly more accurate) to perform the desired (regularized) mode sum in a single step, in a manner that does not require the knowledge of $c_{01}\left(\omega\right)$. To this end, we construct the sequence of partial sums $\sum_{l^{\prime}=0}^{l}\left[T_{uu\left(\omega l^{\prime}\right)}-c_{2}l^{\prime}\left(l^{\prime}+1\right)\right]$ and fit it with $\sim 100$ (or more) orders in $1/\left(l+1\right)$, starting with order $\left(l+1\right)^{1}$. The coefficient of $\left(l+1\right)^{0}$ constitutes the desired regularized mode sum, namely the quantity $T_{uu}^{\text{basic}}\left(\omega\right)$, which is extracted to a precision of $\gtrsim 50$ digits (depending on $r$).
@@ D.1 Analytic extension of the HTPF
where curly brackets denote symmetrization with respect to the spacetime point, that is, for two functions $\zeta$ and $\xi$, we define $\{\zeta(x),\xi(x^{\prime})\}\equiv\zeta(x)\xi(x^{\prime})+\xi(x)\zeta(x^{\prime})$. Concentrating on the polar axis ($\theta=0$, hence only $m=0$ survives), this reduces to
@@
This form corresponds to a purely ingoing wave which may be extended smoothly beyond the EH. However, taking the $r_{*}\to-\infty$ asymptotic behavior of $\psi_{\omega l}^{\text{up}}$ [by setting $m=0$ in Eq. (27)] in the general form of $f_{\omega l}^{\text{up}}$, we find that $f_{\omega l}^{\text{up}}$ has two distinct contributions at the EH: a smooth ingoing $\propto e^{-i\omega v}$ term, and an outgoing $\propto e^{-i\omega u_{\text{ext}}}$ term which is infinitely oscillatory at the EH, as $u_{\text{ext}}\to\infty$ there. (This $\propto e^{-i\omega u_{\text{ext}}}$ term exists in the entire vicinity of $r=r_{+}$, which includes the vicinity of the EH.) The asymptotic form of the “up” modes at the EH vicinity is then 393939 One may notice that the form (144) presented here for the asymptotic behavior of $f_{\omega l}^{\text{up}}$ in the EH limit differs from Eq. (3.15) in Ref. [3] by an additional term proportional to $e^{-i\omega u_{\text{ext}}}$. As elaborated in Ref. [3], Eq. (3.15) therein aligns more closely with the intuitive interpretation of a typical wavepacket originating from $H_{\text{past}}$, then being reflected to the EH and transmitted to FNI (as illustrated in Fig. 2 therein). However, note that while Eq. (3.15) in Ref. [3] conforms with the more intuitive picture commonly accepted in the literature, the form (144) presented here is the exact form (as obtained by simply plugging Eq. (27) into Eq. (29), taking $m=0$). It is this form that facilitates analytic continuation, as described in the current appendix.
@@
On the other side of the EH, within the BH interior, there are the “right” and “left” Eddington modes of Eq. (31) (in which we set $m=0$). Considering the asymptotic behavior of $\psi_{\omega l}^{\text{int}}$ given in Eq. (30), one obtains the asymptotic behavior of $f_{\omega l}^{R}$ and $f_{\omega l}^{L}$ at the EH vicinity:  [^40] The same remark made in footnote 39 regarding the asymptotic behavior of $f^{\text{up}}_{\omega l}$ at the EH is valid here for $f^{L}_{\omega l}$, comparing the above Eq. (146) with Eq. (3.19) of Ref. [3]. As manifested in the latter (see also Fig. 2 therein), the $f^{L}_{\omega l}$ Eddington modes are commonly thought of as arising from $H_{L}$ and having zero initial data at the EH. However, the form given here constitutes the exact asymptotic behavior at $r\to r_{+}$, and hence applies to $H_{L}$ as well as to the EH. While indeed the $f^{L}_{\omega l}$ modes are more naturally tied to $H_{L}$ and may be intuitively thought of as arising from $H_{L}$ (where $u_{\text{int}}$ varies), in the current context of analytical extension we restrict our attention to the asymptotic behavior at the EH vicinity.
@@
Evaluating this function at $z=-|z|$, which corresponds to taking $\phi=\pi$  [^41] Here we analytically extend along a curve in the complex plane bypassing the $r=r_{+}$ singularity from above. A second option would be to bypass the $r=r_{+}$ singularity from below. (Note that since the quantity of interest, the HTPF, is real and analytic across the EH, one may analytically extend along any curve of choice in the complex plane, and the result should be independent of the curve chosen.) This second option would result in replacing $e^{-\alpha\pi}$ in what follows by $e^{\alpha\pi}$. However, the real part of the final mode-sum expression given in Eq. (153) is in fact invariant under this choice of curve. (The imaginary part changes its sign under this change of curve, but this does not concern us, as discussed in footnote 42)., we obtain
@@
The entire $G_{\text{dif}}^{U}\left(x,x^{\prime}\right)$ quantity is expected to vanish in order for the two variants to yield the same results – see footnote 43.

### v2

@@ I Introduction
[…]
@@
The literature results mentioned so far were for spherically-symmetric BHs. In the astrophysically most important case of a stationary, rotating (Kerr) BH, because of the lack of spherical symmetry, $\theta$-splitting is not possible and, because of the lack of staticity, the Euclideanization technique is in principle not implementable either. Thus, progress in Kerr was for a long time made only by calculating differences of renormalized expectation values of quantities in two different states. This was some times combined with the expected behavior of the RSET for one of the states in some specific spacetime region (namely, at infinity or near a horizon) in order to apply the state subtraction method so as to gain knowledge about the behavior of the RSET for the other state in that region: see Refs. [34, 59, 60, 35] outside the EH and Ref. [3] on the CH (see also Ref. [61] extending the latter work, and [4] on the CH of Kerr-dS). An exception to that is the axis of symmetry of Kerr, where no rotation is ‘felt’, and that was used to directly obtain the RSET in a ‘formal’ Hartle-Hawking state on the pole of the EH in Refs. [62, 63]. An important technical breakthrough in renormalization was achieved using the $t$- and $\varphi$-splitting variants of PMR in [64], where the authors managed to calculate the RSET outside the EH of Kerr. This is so far the only time that the calculation of an RSET has been carried out outside a Kerr BH.
@@ II.4 Field quantization and Unruh state
In this subsection we sketch the quantization of the field and the definition of our quantum state of interest, namely the Unruh state. For details, we refer the reader to Secs. III and IV in Ref. [30]. The scalar field $\Phi$ is typically quantized by expanding it in terms of a choice of modes and then promoting the coefficients in the mode expansion to (creation and annihilation) operators. In the previous subsection we introduce the Eddington modes since they are convenient for practical calculations. However, the Unruh state is instead more naturally defined in terms of some other modes, the so-called Unruh modes, which consist of the union of the following two subsets of modes. The first subset, which merges with the Eddington $f^{\text{in}}_{\omega lm}$ in the BH exterior, is defined as having no upward excitations coming from $\{U\in\mathbb{R},V=0\}$ (i.e., from the union of $H_{L}$ and $H_{\text{past}}$, see Fig. 1) and having positive frequencies with respect to the Eddington coordinate $v$ along $\{U\to-\infty,V>0\}$ (i.e., along past null infinity, see Fig. 1); the second subset is defined by being positive frequency with respect to the Kruskal coordinate $U$ along $H_{L}\cup H_{\text{past}}$ and having no upward excitations from past null infinity. The Unruh state [31] is then defined as the quantum state which is annihilated by the annihilation operator coefficients when expanding the quantum field in Unruh modes. The Unruh state is the state of relevance for astrophysical BHs since it models a BH evaporating via the emission of quantum, thermal (Hawking) radiation.
@@ Appendix A Large-$l$ asymptotics
In this Appendix we briefly describe the large-$l$ analysis of the interior radial function $\psi_{\omega l}^{\text{int}}$ [see Eq. (30)] and the reflection coefficient $\rho^{\text{up}}_{\omega l}$ [see Eq. (27)] in a Kerr BH, focusing on the $m=0$ case, given that only $m=0$ modes are relevant at the pole. The analysis we describe yields the leading order expressions which are subsequently used in Sec. III.2.1 for the ID computation of $\left\langle\Phi^{2}\right\rangle$.
@@ B.1 Basic ID analysis
Thus, by taking minus the second derivative with respect to $\varepsilon$ of Eq. (109), we obtain the following distributional identity valid for $\sin\delta\neq\varepsilon$  [^34] In this footnote we present an alternative derivation of Eq. (114). Strictly speaking, this integral is not well defined because the $\omega^{2}$ factor [or the $\omega$ factor in the analogous situation of the integral appearing in Eq. (126)] spoils convergence at large $\omega$. Nevertheless, the integral (like some other integrals involved in our $t$-splitting method) becomes well-defined with the procedure of “oscillation cancellation” (see Ref. [56]), and it is this “generalized integral” that we want to compute here. An alternative (and equivalent) formulation of this generalized integral is via multiplying the integrand by a regulating function $\exp(-c\,\omega)$ (with $c>0$) and then taking the limit $c\to 0$ after integration. This alternative definition, involving the $\exp(-c\,\omega)$ factor, is more commonly used in the mathematical physics literature; however, it is much harder to numerically implement it (compared to oscillation cancellation). Nevertheless, for the analytical derivation of this (generalized) integral – which is our concern here – this second method [i.e. using the regulating function $\exp(-c\,\omega)$] is very convenient. Thus, the relevant integrand now becomes $\omega^{2}\cos\left(\omega\varepsilon\right)J_{0}\left(\omega\sin\delta\right)\exp(-c\,\omega)$. Unfortunately, Mathematica is unable to directly perform this integral. However, when reduced to the evaluation of $\Re\lim_{c\to 0}\int_{0}^{\infty}\omega^{2}\exp\left(i\omega\varepsilon\right)J_{0}\left(\omega\sin\delta\right)\exp(-c\,\omega)$, Mathematica successfully yields the desired result given in Eq. (114).:
@@
Finally, we deal with $P_{\text{div}}^{\left(2\right)}$. We denote
$\Sigma_{2}\left(\delta,\omega\right)\equiv\sum_{l=0}^{\infty}\left(\lambda_{\omega l}\frac{S_{\omega l}\left(\delta\right)}{S_{\omega l}\left(0\right)}\right)\,,$ (117)
@@ Appendix C Numerical methods: the computational side
This section deals with the numerical implementation of the $t$-splitting regularization procedure in practice (which we perform using the software Mathematica [73]) — from the computation of the basic ingredients $\rho_{\omega l}^{\text{up}}$, $\psi_{\omega l}^{\text{int}}$ and $\psi_{\omega l,r}^{\text{int}}$ (as well as $S_{\omega l}(\theta=0)$) to the performance of the various steps (including various challenges that arise), leading to the renormalized quantities $\left\langle\Phi^{2}\right\rangle_{\text{ren}}^{U}$ and $\left\langle T_{yy}\right\rangle_{\text{ren}}^{U}$ in the cases under consideration (detailed in Sec. IV).
@@
Finally, the spheroidal eigenfunctions $S_{\omega l}(\theta=0)$ [defined via Eqs. (44), (20) and (19)] were computed using Mathematica’s built-in function “SpheroidalPS”. In particular, taking normalization into account, our $S_{\omega l}(\theta=0)$ corresponds to $\sqrt{(2l+1)/2}\;\text{SpheroidalPS}[l,0,i\,a\,\omega,1]$. The numerical evaluation of the spheroidal eigenfunctions was done to 300 significant digits.
@@ C.2 Numerical methods
First, in principle, the ID needs to be subtracted from the bare mode contribution, to be followed by a summation over $l$ (per $\omega$) to produce the basic integrand function [defined in Eq. (40)]. For $\Phi^{2}$ this ID subtraction stage (demonstrated in Fig. 3) merely includes the removal of the analytically-known $O\left(l^{0}\right)$ leading order $c_{0}$ [see Eq. (59)], leaving a remainder that decays like $1/l\left(l+1\right)$, which is to be summed over. As already mentioned above, we overcome this slow decay by fitting the sequence of partial sums (namely $\sum_{l^{\prime}=0}^{l}\left(E_{\omega l^{\prime}}-E_{\omega l^{\prime}}^{\text{div}}\right)$) with $\sim 100$ (or more) orders in $1/l$ (starting with order $l^{0}$). 383838 We perform this fit by inputting the last $N$ elements of the sequence and the $N$ coefficients of a power series in $1/l$ (starting from the zeroth order). This results in an algebraic system of $N$ equations and $N$ coefficients, which is solved using Mathematica’s “Solve” function. The value of $N$ considered depends on $r$, but it is typically around 100100~{}100100. In a sample of $r$ and $\omega$ values, we also extracted these coefficients using a least-square fit (using Mathematica’s built-in “LinearModelFit” function), and the results aligned exceptionally well with the explicit computation. The fitted coefficient of this $l^{0}$ term then constitutes the desired $l$-sum. This allows the basic integrand $E^{\text{basic}}\left(\omega\right)$ to be extracted with a precision of typically $\gtrsim 50$ digits, depending on $r$ (this level of precision, reflecting the smallness of the $l$-sum error, was in practice evaluated by comparing fits with different numbers of orders in $1/l$, i.e. varying the number $N$ mentioned in footnote 38).
@@
Hence, we now focus on the direct computation of $\left\langle T_{uu}\right\rangle_{\text{ren}}^{U}$. For $T_{uu}$, the large-$l$ divergent piece $T_{uu\left(\omega l\right)}^{\text{div}}$ has an analytically-known leading-order piece [$c_{2}l\left(l+1\right)$, with $c_{2}$ given in Eq. (74)] and an additional $O\left(l^{0}\right)$ piece [namely $c_{01}\left(\omega\right)$ defined in Eq. (76) ]. This quantity $c_{01}\left(\omega\right)$ is, in principle, not known analytically, but can be extracted numerically. Thus, the straightforward procedure would be to compute the $l$-sum in two stages: first, numerically extracting $c_{01}\left(\omega\right)$, and using it to obtain the regularized sequence $T_{uu\left(\omega l\right)}-T_{uu\left(\omega l\right)}^{\text{div}}$; and then summing this regularized sequence over $l$ to obtain $T_{uu}^{\text{basic}}\left(\omega\right)$. This procedure is demonstrated in Fig. 4. However, we find it simpler and more efficient (and significantly more accurate) to perform the desired (regularized) mode sum in a single step, in a manner that does not require the knowledge of $c_{01}\left(\omega\right)$. To this end, we construct the sequence of partial sums $\sum_{l^{\prime}=0}^{l}\left[T_{uu\left(\omega l^{\prime}\right)}-c_{2}l^{\prime}\left(l^{\prime}+1\right)\right]$ and fit it with $\sim 100$ (or more) orders in $1/\left(l+1\right)$, starting with order $\left(l+1\right)^{1}$ (this fit is performed in a similar fashion to that applied in the $\Phi^{2}$ case described above). The coefficient of $\left(l+1\right)^{0}$ constitutes the desired regularized mode sum, namely the quantity $T_{uu}^{\text{basic}}\left(\omega\right)$, which is extracted to a precision of $\gtrsim 50$ digits (depending on $r$).
@@ D.1 Analytic extension of the HTPF
where, recall, curly brackets denote symmetrization with respect to the spacetime point, that is, for two functions $\zeta$ and $\xi$, we define $\{\zeta(x),\xi(x^{\prime})\}\equiv\zeta(x)\xi(x^{\prime})+\xi(x)\zeta(x^{\prime})$. Concentrating on the polar axis ($\theta=0$, hence only $m=0$ survives), this reduces to
@@
This form corresponds to a purely ingoing wave which may be extended smoothly beyond the EH. However, taking the $r_{*}\to-\infty$ asymptotic behavior of $\psi_{\omega l}^{\text{up}}$ [by setting $m=0$ in Eq. (27)] in the general form of $f_{\omega l}^{\text{up}}$, we find that $f_{\omega l}^{\text{up}}$ has two distinct contributions at the EH: a smooth ingoing $\propto e^{-i\omega v}$ term, and an outgoing $\propto e^{-i\omega u_{\text{ext}}}$ term which is infinitely oscillatory at the EH, as $u_{\text{ext}}\to\infty$ there. (This $\propto e^{-i\omega u_{\text{ext}}}$ term exists in the entire vicinity of $r=r_{+}$, which includes the vicinity of the EH.) The asymptotic form of the “up” modes at the EH vicinity is then 404040 One may notice that the form (144) presented here for the asymptotic behavior of $f_{\omega l}^{\text{up}}$ in the EH limit differs from Eq. (3.15) in Ref. [30] by an additional term proportional to $e^{-i\omega u_{\text{ext}}}$. As elaborated in Ref. [30], Eq. (3.15) therein aligns more closely with the intuitive interpretation of a typical wavepacket originating from $H_{\text{past}}$, then being reflected to the EH and transmitted to future null infinity (as illustrated in Fig. 2 therein). However, note that while Eq. (3.15) in Ref. [30] conforms with the more intuitive picture commonly accepted in the literature, the form (144) presented here is the exact form (as obtained by simply plugging Eq. (27) into Eq. (29), taking $m=0$). It is this form that facilitates analytic continuation, as described in the current appendix.
@@
On the other side of the EH, within the BH interior, there are the “right” and “left” Eddington modes of Eq. (31) (in which we set $m=0$). Considering the asymptotic behavior of $\psi_{\omega l}^{\text{int}}$ given in Eq. (30), one obtains the asymptotic behavior of $f_{\omega l}^{R}$ and $f_{\omega l}^{L}$ at the EH vicinity:  [^41] The same remark made in footnote 40 regarding the asymptotic behavior of $f^{\text{up}}_{\omega l}$ at the EH is valid here for $f^{L}_{\omega l}$, comparing the above Eq. (146) with Eq. (3.19) of Ref. [30]. As manifested in the latter (see also Fig. 2 therein), the $f^{L}_{\omega l}$ Eddington modes are commonly thought of as arising from $H_{L}$ and having zero initial data at the EH. However, the form given here constitutes the exact asymptotic behavior at $r\to r_{+}$, and hence applies to $H_{L}$ as well as to the EH. While indeed the $f^{L}_{\omega l}$ modes are more naturally tied to $H_{L}$ and may be intuitively thought of as arising from $H_{L}$ (where $u_{\text{int}}$ varies), in the current context of analytical extension we restrict our attention to the asymptotic behavior at the EH vicinity.
@@
Evaluating this function at $z=-|z|$, which corresponds to taking $\phi=\pi$  [^42] Here we analytically extend along a curve in the complex plane bypassing the $r=r_{+}$ singularity from above. A second option would be to bypass the $r=r_{+}$ singularity from below. (Note that since the quantity of interest, the HTPF, is real and analytic across the EH, one may analytically extend along any curve of choice in the complex plane, and the result should be independent of the curve chosen.) This second option would result in replacing $e^{-\alpha\pi}$ in what follows by $e^{\alpha\pi}$. However, the real part of the final mode-sum expression given in Eq. (153) is in fact invariant under this choice of curve. (The imaginary part changes its sign under this change of curve, but this does not concern us, as discussed in footnote 43)., we obtain
@@
The entire $G_{\text{dif}}^{U}\left(x,x^{\prime}\right)$ quantity is expected to vanish in order for the two variants to yield the same results – see footnote 44.